EMISSION SPECTROSCOPY AND AB INITIO CALCULATIONS ON TaN

R. S. RAM, Department of Chemistry, University of Arizona, Tucson, AZ 85721; J. LI’EVIN, Université Libre de Bruxelles, Laboratoire de Chimie Physique Moléculaire, CP 160/09, Av. F. D. Roosevelt 50, Bruxelles, Belgium; \underline{P. F. BERNATH}, Department of chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1.